OpenStudy (anonymous):

Solve for y: 3/r-2/y=1/6

OpenStudy (anonymous):

"Solve for y: 3/r - 2/y = 1/6" Put the 2/y on the right side, and the 1/6 on the left. So you've successfully solved for 2/y. Now If u = 2/y, then y = 2/u.